# Gravitational Waves Induced by Scalar Perturbations with a Broken Power-law Peak

### v1

@@ Abstract
We propose an analytical approximation for the energy spectrum of the scalar induced gravitational waves (SIGWs) generated by a broken power-law power spectrum, and find that both the asymptotic power-law tails and the intermediate peak contribute distinct features to the SIGW spectrum. Moreover, the broken power-law power spectrum has abundant near-peak features and our results can be used as a near-peak approximation that covers a wide range of models. Our analytical approximation is useful in the rapid generation of the SIGW energy spectrum, which is beneficial for gravitational wave data analysis.
@@
According to the detection of the Cosmic Microwave Background (CMB) anisotropies Aghanim et al. (2020); Akrami et al. (2020), the amplitude of the scalar power spectrum is constrained to be approximately $A_{s}\approx 2.1\times 10^{-9}$ across scales ranging from megaparsecs (Mpc) to gigaparsecs (Gpc), exhibiting a nearly scale-invariant feature, which is too small to generate detectable SIGWs. However, the information at scales smaller than Mpc lies beyond the reach of CMB detection, leaving open the possibility of significant amplitude of primordial scalar perturbations. In this context, SIGWs can provide constraints on primordial scalar power spectrum, inflationary models, and the expansion history of the universe Ananda et al. (2007); Baumann et al. (2007); Assadullahi and Wands (2010); Inomata and Nakama (2019); Lu et al. (2019); Yuan et al. (2019); Kapadia et al. (2021); Ragavendra (2022); Braglia et al. (2021); Witkowski et al. (2022); Balaji et al. (2022); Alabidi et al. (2013); Inomata et al. (2019a, b); Domènech et al. (2020); Dalianis and Kouvaris (2021); Fu et al. (2020); Kawasaki et al. (2020); Espinosa et al. (2018); Kohri and Terada (2018); Domènech (2020); Unal (2019); Ota (2020); Yuan et al. (2020a); Yuan and Huang (2021a); Zhang et al. (2021); Atal and Domènech (2021); Adshead et al. (2021); Yuan et al. (2023); Ferrante et al. (2023a); Orlofsky et al. (2017); Byrnes et al. (2019); Fu et al. (2020); Kawasaki et al. (2020); Meng et al. (2023).
@@
Moreover, if a large scalar perturbation exceeds a critical value, the overdense region will collapse to form a PBH once the corresponding wavelength reenters the Hubble horizon. In this case, the abundance of PBHs is closely related to the amplitude of the scalar power spectrum, and hence will be tightly constrained by the observations on SIGWs, leading to the significance of studying both PBHs and SIGWs as complementary topics Inomata et al. (2017); Nakama et al. (2017); Garcia-Bellido et al. (2017); Sasaki et al. (2018); Clesse et al. (2018); Inomata et al. (2020); Papanikolaou et al. (2021); Domènech et al. (2021); Franciolini (2021); Cang et al. (2023); Gehrman et al. (2023a); Papanikolaou (2022); Qiu et al. (2023); Escrivà et al. (2022); Meng et al. (2023); Chang et al. (2023); Gehrman et al. (2023b); Ferrante et al. (2023b); Gu et al. (2023); Yuan and Huang (2024); Huang et al. (2024). For reviews on SIGWs, see Yuan and Huang (2021b); Domènech (2021).
@@
Since the first detection of GW by LIGO, numerous GW events have been detected by the LIGO-Virgo-KAGRA collaborations Abbott et al. (2016a, b, 2019, 2021, 2024, 2023). It is possible that some of the events have primordial origins, implying that the corresponding primordial perturbations would generate SIGWs in the frequency band of Pulsar Timing Arrays (PTAs). Recently, hints of the stochastic GW background have been reported by the PTA collaborations Agazie et al. (2023a, b); Afzal et al. (2023); Reardon et al. (2023); Zic et al. (2023); Antoniadis et al. (2023a, b, c); Xu et al. (2023). Alongside various astrophysical and cosmological explanations (see, for example, Bi et al. (2023); Wu et al. (2024)), a potential interpretation is that this signal originates from SIGWs Franciolini et al. (2023); Inomata et al. (2024); Liu et al. (2024); Yi et al. (2023); You et al. (2023); Jin et al. (2023); Balaji et al. (2023); Basilakos et al. (2024a); Liu et al. (2023); Basilakos et al. (2024b). Furthermore, SIGWs are expected to be detected by future space-based GW detectors such as LISA Amaro-Seoane et al. (2017), Taiji Ruan et al. (2020), TianQin Luo et al. (2016) and DECIGO Kawamura et al. (2021). These observations will place constraints on the abundance of PBH potentially addressing the remaining mass window where PBHs could constitute a significant fraction of dark matter Bartolo et al. (2019a, b); Carr et al. (2021).
@@
Based on the semi-analytical formula derived Espinosa et al. (2018); Kohri and Terada (2018); Domènech (2020), the SIGW energy spectrum has been studied for specific cases and some results will be relevant to this work. During the radiation-dominated (RD) era, the infrared (IR) behavior is derived in Yuan et al. (2020a), assuming a general enhanced scalar power spectrum with both IR and ultraviolet (UV) cutoffs. In this case, the GW spectrum scales as $k^{3}\ln^{2}k$ if $k$ is much smaller than the width of remaining interval, while it shows a transition to $k^{2}\ln^{2}k$ if the width is extremely small. As a special case, SIGWs arise from a log-normal scalar power spectrum were studied in Pi and Sasaki (2020) for the widths $\Delta\ll 1$ and $\Delta\gtrsim 1$ respectively. In this paper, we analytically calculate the SIGW spectrum with the scalar power spectrum possessing a broken power-law feature with a smooth transition, which can typically be generated by single-field inflation models Ballesteros and Taoso (2018); Byrnes et al. (2019); Carrilho et al. (2019); Atal and Domènech (2021). This power spectrum also contains abundant features such as near-peak asymmetry and asymptotic power-law tails, which would lead to characteristic structures in the SIGW spectrum.
@@
This paper is organized as follows. We briefly review the formula of GW spectrum during RD era in Sec. II, and consider the broken power-law power spectrum in Sec. III. We present the approximation on the power spectrum first in Sec. III.1 and the GW spectrum is calculated in Sec. III.2, where the asymptotic behavior and the comparison to the numerical result is also presented. Finally, we summarize our findings in Sec. IV. To streamline the calculations, a quick approach can be taken by starting from the final result, Eq. (55), and referring back to the intermediate results as needed.
@@ II SIGW during RD era
In this section, we will briefly review the SIGWs generated during the RD era. Considering scalar and tensor perturbations only, the perturbed Friedmann-Robertson-Walker (FRW) metric in Newtonian gauge can be written as
@@
where $\phi$ is the linear scalar perturbation and $h_{ij}$ refers to the second-order tensor perturbation. For SIGWs in different gauges, see Tomikawa and Kobayashi (2020); De Luca et al. (2020); Inomata and Terada (2020); Yuan et al. (2020b); Giovannini (2020); Lu et al. (2020). During the RD era, the equations of motion for $\phi$ and $h_{ij}$ in Fourier space can be written as
@@
The prime denotes the derivative with respect to the conformal time $\eta$, $\lambda$ denotes the two polarization modes orthogonal to $\bm{k}$, and $\mathcal{S}_{\lambda}$ is the source term up to $\phi_{\bm{k}}^{2}$ order. Here $\phi_{\bm{k}}(\eta)$ and $\mathcal{S}_{\lambda}(\bm{k},\eta)$ take the form:
@@
In the following section, we will show a detailed calculation on Eq. (12) in the circumstance that $\mathcal{P_{R}}$ has a broken power-law feature, which inspires further discussion.
@@ III Broken Power-law Feature
The spectrum has a single peak at $k=k_{\ast}$ and is normalized by $\mathcal{P}_{\mathrm{PL}}(k_{\ast})=A$. After leaving the peak, $\mathcal{P}_{\mathrm{PL}}(k)$ decays as a power-law with the power exponent $\alpha$ and $-\beta$ respectively in the $k\rightarrow 0$ and $k\rightarrow+\infty$ regions. Further discussion on the features of power spectrum will be presented in Sec. III.1, which will be used in the analytical approach to $\Omega_{\mathrm{GW}}$ in Sec. III.2.
The broken power-law peak is typically derived from single-field inflation and the parameter is further restricted by $\alpha<4$ Ballesteros and Taoso (2018); Byrnes et al. (2019); Carrilho et al. (2019); Atal and Domènech (2021). Additionally, the coefficient of near-peak expansion is independent up to the $(k-k_{\ast})^{3}$ order, which means that the broken power-law peak is always a reasonable near-peak approximation even if it is highly asymmetric on both sides. Therefore, we neglect the restriction on the parameter to step further to a general situation.

### v2

@@ Abstract
We give an analytical approximation for the energy spectrum of the scalar-induced gravitational waves (SIGWs) generated by a broken power-law power spectrum, and find that both the asymptotic power-law tails and the intermediate peak contribute distinct features to the SIGW spectrum. Moreover, the broken power-law power spectrum has abundant near-peak features and our results can be used as a near-peak approximation that covers a wide range of models. Our analytical approximation is useful in the rapid generation of the SIGW energy spectrum, which is beneficial for gravitational wave data analysis.
@@
According to the detection of the Cosmic Microwave Background (CMB) anisotropies Aghanim et al. (2020); Akrami et al. (2020), the amplitude of the scalar power spectrum is constrained to be approximately $A_{s}\approx 2.1\times 10^{-9}$ across scales from megaparsecs (Mpc) to gigaparsecs (Gpc), exhibiting a nearly scale-invariant feature, which is too small to generate detectable SIGWs. However, the information at scales smaller than Mpc lies beyond the reach of CMB detection, leaving the possibility of significant enhancement of primordial scalar perturbations. In this context, SIGWs can provide constraints on primordial scalar power spectrum, inflationary models, and the expansion history of the universe Ananda et al. (2007); Baumann et al. (2007); Assadullahi and Wands (2010); Inomata and Nakama (2019); Lu et al. (2019); Yuan et al. (2019); Kapadia et al. (2021); Ragavendra (2022); Braglia et al. (2021); Witkowski et al. (2022); Balaji et al. (2022); Alabidi et al. (2013); Inomata et al. (2019a, b); Domènech et al. (2020); Dalianis and Kouvaris (2021); Fu et al. (2020); Kawasaki et al. (2020); Espinosa et al. (2018); Kohri and Terada (2018); Domènech (2020); Unal (2019); Ota (2020); Yuan et al. (2020a); Yuan and Huang (2021a); Zhang et al. (2021); Atal and Domènech (2021); Adshead et al. (2021); Yuan et al. (2023); Ferrante et al. (2023a); Orlofsky et al. (2017); Byrnes et al. (2019); Fu et al. (2020); Kawasaki et al. (2020); Meng et al. (2023).
@@
Moreover, if a large scalar perturbation exceeds a critical value, the overdense region will collapse to form a PBH once the corresponding wavelength reenters the Hubble horizon. In this case, the abundance of PBHs is closely related to the amplitude of the scalar power spectrum, and hence will be tightly constrained by the observations on SIGWs, leading to the significance of studying both PBHs and SIGWs as complementary topics Inomata et al. (2017); Nakama et al. (2017); Garcia-Bellido et al. (2017); Sasaki et al. (2018); Clesse et al. (2018); Inomata et al. (2020); Papanikolaou et al. (2021); Domènech et al. (2021); Franciolini (2021); Cang et al. (2023); Gehrman et al. (2023a); Papanikolaou (2022); Qiu et al. (2023); Escrivà et al. (2022); Meng et al. (2023); Chang et al. (2023); Gehrman et al. (2023b); Ferrante et al. (2023b); Gu et al. (2023); Yuan and Huang (2024); Huang et al. (2024); Papanikolaou et al. (2024a, b). For reviews on SIGWs, see Yuan and Huang (2021b); Domènech (2021).
@@
Since the first detection of GW by LIGO, numerous GW events have been detected by the LIGO-Virgo-KAGRA collaborations Abbott et al. (2016a, b, 2019, 2021, 2024, 2023). It is possible that some of the events have primordial origins Sasaki et al. (2016); Chen and Huang (2018); Raidal et al. (2019); De Luca et al. (2020a); Hall et al. (2020); Bhagwat et al. (2021); Hütsi et al. (2020); Wong et al. (2021); De Luca et al. (2021); Bavera et al. (2022); Franciolini et al. (2021); Chen et al. (2022); Chen and Hall (2024); Yuan and Huang (2024); Huang et al. (2024), implying that the corresponding primordial perturbations would generate SIGWs in the frequency band of Pulsar Timing Arrays (PTAs) (see e.g., Yuan and Huang (2024)). Recently, hints of the stochastic GW background have been reported by the PTA collaborations Agazie et al. (2023a, b); Afzal et al. (2023); Reardon et al. (2023); Zic et al. (2023); Antoniadis et al. (2023a, b, c); Xu et al. (2023). Alongside various astrophysical and cosmological explanations (see, for example, Bi et al. (2023); Wu et al. (2024)), a potential interpretation is that this signal originates from SIGWs Franciolini et al. (2023); Inomata et al. (2024); Liu et al. (2024); Yi et al. (2023); You et al. (2023); Jin et al. (2023); Balaji et al. (2023); Basilakos et al. (2024a); Liu et al. (2023); Basilakos et al. (2024b). Furthermore, SIGWs are expected to be detected by future space-based GW detectors such as LISA Amaro-Seoane et al. (2017), Taiji Ruan et al. (2020), TianQin Luo et al. (2016) and DECIGO Kawamura et al. (2021). These observations will place constraints on the abundance of PBH potentially addressing the remaining mass window where PBHs could constitute a significant fraction of dark matter Bartolo et al. (2019a, b); Carr et al. (2021).
@@
Based on the semi-analytical formula derived in Espinosa et al. (2018); Kohri and Terada (2018); Domènech (2020), the SIGW energy spectrum has been studied for specific cases and some results will be relevant to this work. During the radiation-dominated (RD) era, the infrared (IR) behavior is derived in Yuan et al. (2020a), assuming a general enhanced scalar power spectrum with both IR and ultraviolet (UV) cutoffs. In this case, the GW spectrum scales as $k^{3}\ln^{2}k$ if $k$ is much smaller than the width of remaining interval, while it shows a transition to $k^{2}\ln^{2}k$ if the width is extremely small. As a special case, SIGWs arise from a log-normal scalar power spectrum were studied in Pi and Sasaki (2020) for the widths $\Delta\ll 1$ and $\Delta\gtrsim 1$ respectively. In this paper, we analytically calculate the SIGW spectrum with the scalar power spectrum possessing a broken power-law feature with a smooth transition, which can typically be generated by single-field inflation models Ballesteros and Taoso (2018); Byrnes et al. (2019); Carrilho et al. (2019); Atal and Domènech (2021). This power spectrum also contains abundant features such as near-peak asymmetry and asymptotic power-law tails, which would lead to characteristic structures in the SIGW spectrum.
@@
This paper is organized as follows. We briefly review the formula of SIGW spectrum during RD in Sec. II, and consider the broken power-law power spectrum in Sec. III. In Sec. III.1, we give the analytical result of SIGW spectrum generated by broken power-law power spectrum, and compare with the numerical result. In Sec. III.2 and Sec. III.3, we give the IR and UV behavior of our result. Finally, we summarize our findings in Sec. IV. To streamline the calculations, a quick approach can be taken by starting from the final result, Eq. (30), and referring back to the intermediate results as needed.
@@ II SIGW during RD era
In this section, we will briefly review the SIGWs generated during the RD era. Consider only scalar and tensor perturbations, the perturbed Friedmann-Robertson-Walker (FRW) metric in Newtonian gauge can be written as
@@
where $\phi$ is the linear scalar perturbation and $h_{ij}$ refers to the second-order tensor perturbation. For SIGWs in different gauges, see Tomikawa and Kobayashi (2020); De Luca et al. (2020b); Inomata and Terada (2020); Yuan et al. (2020b); Giovannini (2020); Lu et al. (2020). During the RD era, the equations of motion for $\phi$ and $h_{ij}$ in Fourier space can be written as
@@
The prime denotes the derivative with respect to the conformal time $\eta$. $\lambda$ denotes the two polarization modes of GWs, and $\mathcal{S}_{\lambda}$ is the source term up to $\phi_{\bm{k}}^{2}$ order. Here $\phi_{\bm{k}}(\eta)$ and $\mathcal{S}_{\lambda}(\bm{k},\eta)$ take the form:
@@
In the following section, we will show a detailed calculation based on Eq. (12) in the circumstance that $\mathcal{P_{R}}$ has a broken power-law feature.